OpenStudy (safa102):

What is the slope-intercept form of the function that contains the points (6,2) and (4,8)?

OpenStudy (safa102):

not really, that why i put it up!

OpenStudy (safa102):

?

OpenStudy (safa102):

2 - 8 = 6 6 - 4 = 2

OpenStudy (safa102):

oh ok

OpenStudy (safa102):

m = -3 b = 2

OpenStudy (safa102):

i think?

OpenStudy (safa102):

ok that's what i put in right

OpenStudy (safa102):

final answer

OpenStudy (anonymous):

\[\frac{2-8}{6-4}\] \[\frac{-6}{2}\] m = -3 y = mx + b (now choose any of the points, either (6,2) or (4,8) Using (6,2) : x=6 y=2 (2) = (-3)(6) + b 2 = -18 + b (we now move 18 to the other side and change it from negative to positive) 2 +18 = b 20 = b (b = y-intercept) Using (4,8): x=4 y=8 (8) = (-3)(4) + b 8 = -12 + b 8 + 12 = b 20 = b *using either point will get you the same answer.

OpenStudy (safa102):

thanks so much

OpenStudy (anonymous):

np